# Orbital optimized vs time-dependent density functional calculations of intramolecular charge transfer excited states

### v3

@@ 3.0.2 Excitation energy
As the extent of charge transfer increases, TD-DFT more significantly underestimates the excitation energy, yielding absolute errors up to eV for some of the excitations with longest charge transfer distance. As a consequence, TD-DFT predicts the A states of planar PP with strong charge transfer character to be lower in energy than the more local B state, which is in disagrement with high-level quantum chemistry calculations (see Table 1). In contrast, the errors of the orbital optimized calculations are always within eV, and no negative slope with increasing charge transfer distance is observed. As a result, the order of the A and B states of planar PP is predicted correctly. These trends are the same for the LDA and the BLYP functional, as shown in the Supporting Information. A similar correlation between the underestimation of the excitation energy and the extent of charge transfer, though less pronounced, has been observed previously for TD-DFT calculations with the global hybrid functionals B3LYP and PBE015, 9, which include 20% and 25% of exact exchange, respectively. For the excitations with charge transfer distance bigger than 2.0 Å, the orbital optimized calculations with the LDA and GGA functionals (MAE 0.5 eV) achieve better accuracy than TD-DFT with the more computationally expensive global hybrid functionals (MAE of 0.73 eV and 0.91 eV for B3LYP and PBE0, respectively9).
@@
Using the global hybrid functional B3LYP reduces the errors of both the orbital optimized and TD-DFT calculations. However, for the excitations with long-range charge transfer, the absolute errors of TD-DFT remain large, up to eV, while the absolute errors of the orbital optimized calculations are less than 0.6 eV. For the B state of twisted DMABN, the errors of the TD-DFT calculations with the BLYP and B3LYP functionals are bigger than for the A state of NDMAN, despite the latter having a longer charge transfer distance. This is because the spatial overlap between the orbitals involved in the B excitation of twisted DMABN is significantly smaller than for the A state of NDMAN. As expected16, 9, 19, the TD-DFT calculations with the range-separated hybrid functional CAM-B3LYP provide small errors for the excited states with long-range charge transfer but overestimate the excitation energy of the excited state with short charge transfer distance. An improvement of the excitation energy of long-range charge transfer states is also observed for the orbital optimized calculations with CAM-B3LYP compared to BLYP and B3LYP. There, the errors are bigger than for TD-DFT, but still relatively small (-0.15 eV). Moreover, the orbital optimized calculations with CAM-B3LYP overestimate the excitation energy of the two short-range charge transfer states slightly less than TD-DFT.
@@ 4 Discussion and conclusions
State specific orbital optimization in time-independent density functional calculations of excited states offers a distinct advantage over TD-DFT, particularly for describing high-energy and charge transfer excitations. However, orbital optimized calculations face two main challenges: (1) The excited state solutions are typically saddle points on the electronic energy surface, and (2) the non-linearity of the orbital optimization leads to a large number of stationary solutions, some lacking a clear association with a specific electronic state. Here, it is found that the presence of charge-delocalized solutions affects calculations of intramolecular charge transfer states with both local, semi-local, and hybrid functionals. Exploratory orbital optimized calculations with the B3LYP functional and using a DIIS-based approach have been performed for the A excited state of the twisted PP molecule where a charge-localized and a charge-delocalized solution are found for the local and semi-local functionals (see section 2.3). Similarly to the DO-MOM calculations with the PBE functional illustrated in section 2.3, the B3LYP calculations converge to a mixed-spin solution where mixing between the hole initially localized on the pyrrole group and an occupied orbital initially localized on the phenyl group of the molecule leads to charge delocalization. This solution has a dipole moment and charge transfer distance of 3.82 D and 2.06 Å, respectively, much smaller than those of the charge-localized solutions obtained for the calculations with local and semi-local functionals (7.3-9.5 D and 2.2-2.4 Å, respectively), the latter being in agreement with the TD-DFT calculations using the CAM-B3LYP functionals (dipole moment and charge transfer distance of 9.56 D and 2.35 Å, respectively).
@@
The issue of multiple orbital optimized solutions has been observed for other classes of excitations. For example, it has recently been shown that the presence of charge-delocalized solutions affects orbital optimized density functional calculations of core excitations in symmetric molecules65, such as CO and N, and charge transfer states of systems, such as HO-Li, where the donor and acceptor are separated by a large distance, even when global or range-separated hybrid functionals are used17. Convergence on such unphysical solutions can have detrimental effects on the calculation of excited state potential energy surfaces and molecular dynamics simulations, therefore care must be taken to avoid converging on spurious solutions where the charge is too delocalized. Here, a freeze-and-release DO approach involving a first constrained optimization step where the orbitals of the hole and the excited electron are frozen has been presented and shown to be highly effective in converging on charge-localized solutions when applied in calculations of intramolecular charge transfer excited states of organic molecules with the LDA and GGA functionals. Significantly, in the case of the A state of benzonitrile, the charge-localized solution could be obtained even if the electronic distance from the initial guess consisting of ground state canonical orbitals with non-aufbau occupation is larger than the distance to a charge-delocalized solution ( of 0.49 and 0.19, respectively, when using the PBE functional). An extensive assessment of the robustness and efficiency of the method in comparison with other algorithms for orbital optimized excited state calculations, such as MOM, and the application to hybrid functionals will be presented in future works.
@@
The freeze-and-release DO strategy has enabled a benchmark study of the performance of orbital optimized density functional calculations with local and semi-local functionals for a large set of charge transfer excitations in organic molecules. Orbital optimized calculations with the LDA and the PBE and BLYP functionals are found to provide on average better estimates of the excitation energy compared to linear-response TD-DFT calculations with the same functionals. Most importantly, the TD-DFT errors tend to increase with the charge transfer distance. As a consequence, in cases where a molecule has excited states with different characters, TD-DFT can produce an incorrect order of the states because the energy of states with strong charge transfer is underestimated more than those with weak charge transfer, as exemplified by the B and A states of planar PP. In contrast, for the orbital optimized approach, excitations with strong charge transfer character are described with similar accuracy as excitations with a more local character. For the excitations with strong charge transfer, the accuracy of the orbital optimized calculations with the LDA, and the PBE and BLYP functionals rivals the accuracy of TD-DFT with more computationally demanding hybrid functionals9. When comparing charge transfer distances, the deviation between the results of orbital optimized and TD-DFT calculations is small, even for the states with strong charge transfer character, where TD-DFT with semi-local functionals gives large errors on the excitation energy. However, a few limitations of the TD-DFT calculations should be noted. Firstly, even when the full, relaxed TD-DFT difference density matrix is used, the charge transfer distance obtained when using local and semi-local functionals tends to be large in comparison to results obtained in orbital optimized calculations. Therefore, linear-response TD-DFT seems to lack relaxation effects compared to the orbital optimized calculations, which are fully variational. Secondly, the charge transfer distance computed with TD-DFT can be too low in some cases because of mixing with Rydberg states when using diffuse functions, an issue that does not affect the orbital optimized calculations.
@@
The balanced treatment of several electronic states using computationally efficient functionals combined with the robustness of DO can be especially useful for large molecular dynamics simulations in the condensed phase, where lowering the computational cost is crucial for obtaining statistically significant results, but previous orbital optimized approaches have suffered from convergence issues32. Despite these advantages, the absolute errors obtained here for local and semi-local functionals for the excitation energy can still be relatively large (up to eV), while many applications require a higher accuracy. Orbital optimized calculations of excited states can be affected by the delocalization error inherent in local and semi-local KS functionals, which can lead to the errors on the excitation energy observed here. The issue is evident in the case of the A excited state of PP. Previous high-level calculations show minor changes in the character of the excitation with variations in molecular geometry9. However, the solution obtained from orbital optimized calculations at the planar geometry has a relatively high degree of charge delocalization and no charge-localized solution closer to the initial guess has been found, compared to the solution obtained for the same state at the twisted geometry. The presence of charge-delocalized solutions is found to affect both the calculations of the spin-mixed and triplet excited states. Here, state specific methods accounting for the multi-determinant nature of open-shell singlet states in the orbital optimization, such as the restricted open-shell KS approach103, 104, might have an advantage over single-determinant methods, as they avoid calculating both a spin-mixed and a triplet solution.

### v4

@@ 3.0.2 Excitation energy
As the extent of charge transfer increases, TD-DFT more significantly underestimates the excitation energy, yielding absolute errors up to eV for some of the excitations with longest charge transfer distance. As a consequence, TD-DFT predicts the A1 states of planar PP with strong charge transfer character to be lower in energy than the more local B2 state, which is in disagrement with high-level quantum chemistry calculations (see Table 1). In contrast, the errors of the orbital optimized calculations are always within eV, and no negative slope with increasing charge transfer distance is observed. As a result, the order of the A1 and B2 states of planar PP is predicted correctly. These trends are the same for the LDA and the BLYP functional, as shown in the Supporting Information. A similar correlation between the underestimation of the excitation energy and the extent of charge transfer, though less pronounced, has been observed previously for TD-DFT calculations with the global hybrid functionals B3LYP and PBE015, 9, which include 20% and 25% of exact exchange, respectively. For the excitations with charge transfer distance bigger than 2.0 Å, the orbital optimized calculations with the LDA and GGA functionals (MAE 0.5 eV) achieve better accuracy than TD-DFT with the more computationally expensive global hybrid functionals (MAE of 0.73 eV and 0.91 eV for B3LYP and PBE0, respectively9).
@@
Using the global hybrid functional B3LYP reduces the errors of both the orbital optimized and TD-DFT calculations. However, for the excitations with long-range charge transfer, the absolute errors of TD-DFT remain large, up to eV, while the absolute errors of the orbital optimized calculations are less than 0.6 eV. For the B1 state of twisted DMABN, the errors of the TD-DFT calculations with the BLYP and B3LYP functionals are bigger than for the A1 state of NDMAN, despite the latter having a longer charge transfer distance. This is because the spatial overlap between the orbitals involved in the B1 excitation of twisted DMABN is significantly smaller than for the A1 state of NDMAN. As expected16, 9, 19, the TD-DFT calculations with the range-separated hybrid functional CAM-B3LYP provide small errors for the excited states with long-range charge transfer but overestimate the excitation energy of the excited state with short charge transfer distance. An improvement of the excitation energy of long-range charge transfer states is also observed for the orbital optimized calculations with CAM-B3LYP compared to BLYP and B3LYP. There, the errors are bigger than for TD-DFT, but still relatively small (-0.15 eV). Moreover, the orbital optimized calculations with CAM-B3LYP overestimate the excitation energy of the two short-range charge transfer states slightly less than TD-DFT.
@@ 4 Discussion and conclusions
State specific orbital optimization in time-independent density functional calculations of excited states offers a distinct advantage over TD-DFT, particularly for describing high-energy and charge transfer excitations. However, orbital optimized calculations face two main challenges: (1) The excited state solutions are typically saddle points on the electronic energy surface, and (2) the non-linearity of the orbital optimization leads to a large number of stationary solutions, some lacking a clear association with a specific electronic state. Here, it is found that the presence of charge-delocalized solutions affects calculations of intramolecular charge transfer states with both local, semi-local, and hybrid functionals. Exploratory orbital optimized calculations with the B3LYP functional and using a DIIS-based approach have been performed for the A1 excited state of the twisted PP molecule where a charge-localized and a charge-delocalized solution are found for the local and semi-local functionals (see section 2.3). Similarly to the DO-MOM calculations with the PBE functional illustrated in section 2.3, the B3LYP calculations converge to a mixed-spin solution where mixing between the hole initially localized on the pyrrole group and an occupied orbital initially localized on the phenyl group of the molecule leads to charge delocalization. This solution has a dipole moment and charge transfer distance of 3.82 D and 2.06 Å, respectively, much smaller than those of the charge-localized solutions obtained for the calculations with local and semi-local functionals (7.3-9.5 D and 2.2-2.4 Å, respectively), the latter being in agreement with the TD-DFT calculations using the CAM-B3LYP functionals (dipole moment and charge transfer distance of 9.56 D and 2.35 Å, respectively).
@@
The issue of multiple orbital optimized solutions has been observed for other classes of excitations. For example, it has recently been shown that the presence of charge-delocalized solutions affects orbital optimized density functional calculations of core excitations in symmetric molecules65, such as CO2 and N2, and charge transfer states of systems, such as H2O-Li+, where the donor and acceptor are separated by a large distance, even when global or range-separated hybrid functionals are used17. Convergence on such unphysical solutions can have detrimental effects on the calculation of excited state potential energy surfaces and molecular dynamics simulations, therefore care must be taken to avoid converging on spurious solutions where the charge is too delocalized. Here, a freeze-and-release DO approach involving a first constrained optimization step where the orbitals of the hole and the excited electron are frozen has been presented and shown to be highly effective in converging on charge-localized solutions when applied in calculations of intramolecular charge transfer excited states of organic molecules with the LDA and GGA functionals. Significantly, in the case of the A2 state of benzonitrile, the charge-localized solution could be obtained even if the electronic distance from the initial guess consisting of ground state canonical orbitals with non-aufbau occupation is larger than the distance to a charge-delocalized solution ( of 0.49 and 0.19, respectively, when using the PBE functional). An extensive assessment of the robustness and efficiency of the method in comparison with other algorithms for orbital optimized excited state calculations, such as MOM, and the application to hybrid functionals will be presented in future works.
@@
The freeze-and-release DO strategy has enabled a benchmark study of the performance of orbital optimized density functional calculations with local and semi-local functionals for a large set of charge transfer excitations in organic molecules. Orbital optimized calculations with the LDA and the PBE and BLYP functionals are found to provide on average better estimates of the excitation energy compared to linear-response TD-DFT calculations with the same functionals. Most importantly, the TD-DFT errors tend to increase with the charge transfer distance. As a consequence, in cases where a molecule has excited states with different characters, TD-DFT can produce an incorrect order of the states because the energy of states with strong charge transfer is underestimated more than those with weak charge transfer, as exemplified by the B2 and A1 states of planar PP. In contrast, for the orbital optimized approach, excitations with strong charge transfer character are described with similar accuracy as excitations with a more local character. For the excitations with strong charge transfer, the accuracy of the orbital optimized calculations with the LDA, and the PBE and BLYP functionals rivals the accuracy of TD-DFT with more computationally demanding hybrid functionals9. When comparing charge transfer distances, the deviation between the results of orbital optimized and TD-DFT calculations is small, even for the states with strong charge transfer character, where TD-DFT with semi-local functionals gives large errors on the excitation energy. However, a few limitations of the TD-DFT calculations should be noted. Firstly, even when the full, relaxed TD-DFT difference density matrix is used, the charge transfer distance obtained when using local and semi-local functionals tends to be large in comparison to results obtained in orbital optimized calculations. Therefore, linear-response TD-DFT seems to lack relaxation effects compared to the orbital optimized calculations, which are fully variational. Secondly, the charge transfer distance computed with TD-DFT can be too low in some cases because of mixing with Rydberg states when using diffuse functions, an issue that does not affect the orbital optimized calculations.
@@
The balanced treatment of several electronic states using computationally efficient functionals combined with the robustness of DO can be especially useful for large molecular dynamics simulations in the condensed phase, where lowering the computational cost is crucial for obtaining statistically significant results, but previous orbital optimized approaches have suffered from convergence issues32. Despite these advantages, the absolute errors obtained here for local and semi-local functionals for the excitation energy can still be relatively large (up to eV), while many applications require a higher accuracy. Orbital optimized calculations of excited states can be affected by the delocalization error inherent in local and semi-local KS functionals, which can lead to the errors on the excitation energy observed here. The issue is evident in the case of the A1 excited state of PP. Previous high-level calculations show minor changes in the character of the excitation with variations in molecular geometry9. However, the solution obtained from orbital optimized calculations at the planar geometry has a relatively high degree of charge delocalization and no charge-localized solution closer to the initial guess has been found, compared to the solution obtained for the same state at the twisted geometry. The presence of charge-delocalized solutions is found to affect both the calculations of the spin-mixed and triplet excited states. Here, state specific methods accounting for the multi-determinant nature of open-shell singlet states in the orbital optimization, such as the restricted open-shell KS approach103, 104, might have an advantage over single-determinant methods, as they avoid calculating both a spin-mixed and a triplet solution.